If a gambler rolls two dice and gets the sum of 7, he wins $20. If he gets a sum of 4, he wins $40. The cost of playing is $14. What is the expectation of this game?
Answer by Boreal(15235)

P(7)=1/6. It pays $20, the product is 20/6 or $3.33
p(4)=1/12. It pays $40, the product is 40/12 or 20/6, or $3.33
p of everything else is 1-(1/6+1/12)=3/4. It cost $14, so that product is $10.50
$6.67-$10.50=minus $3.83 per time played.